# Nearly-Optimal Bandit Learning in Stackelberg Games with Side Information

### v1

@@ Appendix B Appendix for Section 4: Other Applications
$\pi^{*}(\mathbf{z})=\arg\max_{\mathbf{b}\in[0,1]^{m}}\mathbb{E}_{\boldsymbol{\theta}\sim\mathcal{P}}[u(\mathbf{z},\mathbf{b},\boldsymbol{\theta})]=\arg\max_{\mathbf{b}\in[0,1]^{m}}\sum_{i=1}^{K}\mathbf{p}[i]\cdot u(\mathbf{z},\mathbf{b},\boldsymbol{\theta}^{(i)}),$
where $\mathcal{P}$ is an unknown distribution with support on $\{\boldsymbol{\theta}^{(1)},\ldots,\boldsymbol{\theta}^{(K)}\}$. Let $\pi^{\prime}:=\pi^{(\boldsymbol{\omega})}$ be the optimal policy in the discretization and let $\mathcal{P}^{\prime}$ be the corresponding distribution over $\boldsymbol{\theta}$. We have that
$R(T)$ $=\mathbb{E}_{\boldsymbol{\theta}\sim\mathcal{P}}[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\boldsymbol{\theta})-u(\mathbf{z}_{t},\mathbf{b}_{t},\boldsymbol{\theta})]$
$=\sum_{t=1}^{T}\left(\sum_{i=1}^{K}\mathbf{p}^{\prime}[i](u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\boldsymbol{\theta}^{(i)})-u(\mathbf{z}_{t},\mathbf{b}_{t},\boldsymbol{\theta}^{(i)}))+\sum_{i=1}^{K}(\mathbf{p}^{\prime}[i]-\mathbf{p}[i])\cdot u(\mathbf{z}_{t},\mathbf{b}_{t},\boldsymbol{\theta}^{(i)})\right.$
$+\left.\sum_{i=1}^{K}(\mathbf{p}[i]-\mathbf{p}^{\prime}[i])\cdot u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\boldsymbol{\theta}^{(i)})+\sum_{i=1}^{K}\mathbf{p}^{\prime}[i](u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\boldsymbol{\theta}^{(i)})-u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\boldsymbol{\theta}^{(i)}))\right)$
$\leq 2K+\mathbb{E}_{\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{T}\sim\mathcal{P}^{\prime}}\left[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\boldsymbol{\theta}_{t})-u(\mathbf{z}_{t},\mathbf{b}_{t},\boldsymbol{\theta}_{t})\right]$
@@
The proof follows identically to the proof of Theorem 3.2, but without the loss in utility due to discretization. To see why, let $N_{i}:=\sum_{t:\boldsymbol{\theta}_{t}=\boldsymbol{\theta}^{(i)}}1$ and observe that
$\pi^{*}(\mathbf{z})$ $:=\arg\max_{\mathbf{b}\in[0,1]^{m}}\mathbb{E}_{\mathbf{z}\sim\mathcal{P}}[\sum_{t=1}^{T}u(\mathbf{z},\mathbf{b},\boldsymbol{\theta}_{t})]$
$=\arg\max_{\mathbf{b}\in[0,1]^{m}}\sum_{i=1}^{K}\frac{N_{i}}{T}\cdot\mathbb{E}_{\mathbf{z}\sim\mathcal{P}}[u(\mathbf{z},\mathbf{b},\boldsymbol{\theta}^{(i)})]$

### v2

@@ Appendix B Appendix for Section 4: Other applications
Input: Linear contextual bandit algorithm $\mathcal{R}$
for $t=1,\ldots,T$ do
   Observe $\mathbf{z}_{t}$, compute utility set $U_{t}$
   Let $\mathbf{v}_{t}\leftarrow\mathcal{R}.\mathrm{recommend}(U_{t})$
   Play the action which induces $\mathbf{v}_{t}$
   Receive utility $u_{t}$ and call $\mathcal{R}.\mathrm{observeUtility}(\mathbf{v}_{t},u_{t})$
   end for
Algorithm 3 Reduction for Auctions and Persuasion
@@
$\pi^{*}(\mathbf{z})=\arg\max_{\mathbf{b}\in[0,1]^{m}}\mathbb{E}_{\bm{\theta}\sim\mathcal{P}}[u(\mathbf{z},\mathbf{b},\bm{\theta})]=\arg\max_{\mathbf{b}\in[0,1]^{m}}\sum_{i=1}^{K}\mathbf{p}[i]\cdot u(\mathbf{z},\mathbf{b},\bm{\theta}^{(i)}),$
where $\mathcal{P}$ is an unknown distribution with support on $\{\bm{\theta}^{(1)},\ldots,\bm{\theta}^{(K)}\}$. Let $\pi^{\prime}:=\pi^{(\bm{\omega})}$ be the optimal policy in the discretization and let $\mathcal{P}^{\prime}$ be the corresponding distribution over $\bm{\theta}$. We have that
$R(T)$ $=\mathbb{E}_{\bm{\theta}\sim\mathcal{P}}[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\bm{\theta})-u(\mathbf{z}_{t},\mathbf{b}_{t},\bm{\theta})]$
$=\sum_{t=1}^{T}\left(\sum_{i=1}^{K}\mathbf{p}^{\prime}[i](u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\bm{\theta}^{(i)})-u(\mathbf{z}_{t},\mathbf{b}_{t},\bm{\theta}^{(i)}))+\sum_{i=1}^{K}(\mathbf{p}^{\prime}[i]-\mathbf{p}[i])\cdot u(\mathbf{z}_{t},\mathbf{b}_{t},\bm{\theta}^{(i)})\right$.
$+\left.\sum_{i=1}^{K}(\mathbf{p}[i]-\mathbf{p}^{\prime}[i])\cdot u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\bm{\theta}^{(i)})+\sum_{i=1}^{K}\mathbf{p}^{\prime}[i](u(\mathbf{z}_{t},\pi^{*}(\mathbf{z}_{t}),\bm{\theta}^{(i)})-u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\bm{\theta}^{(i)}))\right)$
$\leq 2K+\mathbb{E}_{\bm{\theta}_{1},\ldots,\bm{\theta}_{T}\sim\mathcal{P}^{\prime}}\left[\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi^{\prime}(\mathbf{z}_{t}),\bm{\theta}_{t})-u(\mathbf{z}_{t},\mathbf{b}_{t},\bm{\theta}_{t})\right]$
@@
The proof follows identically to the proof of Theorem 3.2, but without the loss in utility due to discretization. To see why, let $N_{i}:=\sum_{t:\bm{\theta}_{t}=\bm{\theta}^{(i)}}1$ and observe that
$\pi^{*}(\mathbf{z})$ $=\arg\max_{\mathbf{b}\in[0,1]^{m}}\mathbb{E}_{\mathbf{z}\sim\mathcal{P}}[\sum_{t=1}^{T}u(\mathbf{z},\mathbf{b},\bm{\theta}_{t})]$
$=\arg\max_{\mathbf{b}\in[0,1]^{m}}\sum_{i=1}^{K}\frac{N_{i}}{T}\cdot\mathbb{E}_{\mathbf{z}\sim\mathcal{P}}[u(\mathbf{z},\mathbf{b},\bm{\theta}^{(i)})]$